I can't quite figure out how to create an equation for this word problem
A university bookstore recently sold a wirebound graph-paper notebook for $2.50, and a college-ruled notebook for $2.30. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $118.60. How many of each type were sold?

Answer by scott8148(6628)

let x=graph-paper, so 50-x=college-ruled

x(2.50)+(50-x)(2.30)=118.60 __ multiplying by 10 to clear decimals __ 25x+1150-23x=1186

2x=36 __ x=18